### v1

# Black hole interiors in the thermal view of scalar-tensor gravity

Abstract
The thermal view of scalar-tensor gravity is an analogy with a dissipative fluid. The scalar degree of freedom excites gravity to a positive “temperature”, while Einstein gravity is the “zero-temperature” equilibrium state. We extend this thermal analogy to the interior region near the singularity of spherical, vacuum, uncharged black holes by using the universality of the Kasner behaviour near spacelike singularities. The singularity is “hot”, meaning that gravity diverges from general relativity. The discussion is then extended to black holes in the presence of perfect or imperfect fluids with constant equation of state—the latter determines whether Einstein gravity is approached or not.
Introduction—In spite of its success, the standard theory of gravity, General Relativity (GR) has shortcomings. It predicts the divergence of physical and geometrical quantities and the failure of the spacetime structure itself at spacetime singularities inside black holes and in the early universe. These singularities are ultimately expected to be cured by quantum gravity. In any case, the lowest-order quantum corrections to GR “break” it by introducing extra degrees of freedom or higher order field equations (e.g., [1, 2, 3, 4, 5]). Furthermore, the standard cosmological model based on GR, the $\Lambda$-Cold Dark Matter model, is now subject to various tensions [6, 7]. Finally, the current acceleration of the cosmic expansion discovered in 1998 with Type Ia supernovae was explained with a completely ad hoc dark energy of unknown origin [8], which is far from satisfactory. Cosmologists often prefer the alternative of modifying gravity at cosmological scales without invoking dark energy. For this purpose, the most well studied and popular alternative to GR is $f(R)$ gravity, where the Lagrangian density $f$ is a non-linear function of the Ricci scalar $R$ [9, 10, 11].
The simplest modification of GR consists of adding a scalar degree of freedom $\phi$, as in scalar-tensor (ST) gravity, which corresponds to the inverse of the effective gravitational coupling strength $G_{\mathrm{eff}}\simeq 1/\phi$—the latter becomes a dynamical field sourced by matter already in Brans-Dicke gravity [12], later generalized to what are now called “first-generation” ST theories [13, 14, 15, 16]. The class of $f(R)$ theories popular to explain the cosmic acceleration without dark energy is a subclass of ST gravity. ST gravity with field equations of order not higher than two has been further generalized with the rediscovery of Horndeski gravity [17] and its further DHOST extension, which are now the subject of a large literature.
First-generation ST gravity [12, 13, 14, 15, 16] is described by the Jordan frame action
$S_{\mathrm{ST}}=\int d^{4}x\,\frac{\sqrt{-g}}{16\pi}\left[\phi R-\frac{\omega(\phi)}{\phi}\,\nabla^{c}\phi\nabla_{c}\phi-V(\phi)\right]+S^{\mathrm{(m)}}\,,$ (1)
where (following the notation of Ref. [18] and using units in which the speed of light $c$ and Newton’s constant $G$ are unity) $g_{ab}$ is the spacetime metric with determinant $g$, covariant derivative operator $\nabla_{a}$, and Ricci scalar $R$, $\omega$ is the “Brans-Dicke coupling” (restricted to values $\omega>-3/2$ to avoid the scalar $\phi$ having phantom nature and being subject to fatal instabilities), while $S^{\mathrm{(m)}}$ is the matter action. The Jordan frame field equations derived from this action are
$R_{ab}-\frac{1}{2}\,g_{ab}R=\frac{8\pi}{\phi}\,T_{ab}^{\mathrm{(m)}}+T_{ab}^{(\phi)}\,,$ (2)
$\,\,\,T_{ab}^{(\phi)}\equiv\frac{\omega}{\phi^{2}}\left(\nabla_{a}\phi\nabla_{b}\phi-\frac{1}{2}\,g_{ab}\nabla^{c}\phi\nabla_{c}\phi\right)$
$\qquad\quad\quad+\frac{1}{\phi}\left(\nabla_{a}\nabla_{b}\phi-g_{ab}\Box\phi\right)-\frac{V}{2\phi}g_{ab}\,,\,$
$(2\omega+3)\Box\phi=8\pi T^{\mathrm{(m)}}+\phi\,V^{\prime}-2V-\omega^{\prime}\,\nabla^{c}\phi\nabla_{c}\phi\,,$
where $R_{ab}$ is the Ricci tensor, $T_{ab}^{\mathrm{(m)}}$ is the matter stress-energy tensor with trace $T^{(m)}\equiv g^{ab}T_{ab}^{\mathrm{(m)}}$, $\Box\equiv g^{ab}\nabla_{a}\nabla_{b}$, and a prime denotes differentiation with respect to $\phi$. $T_{ab}^{(\phi)}$ is an effective stress-energy tensor of $\phi$, which originates the thermal view of ST gravity [19, 20, 21, 22, 23, 24]. Provided that the scalar field gradient $\nabla^{c}\phi$ is timelike and future-oriented, it can be used to define the 4-velocity of an effective $\phi$-fluid as
$u^{a}\equiv\frac{\nabla^{a}\phi}{\sqrt{-\nabla^{c}\phi\nabla_{c}\phi}}\,,$ (4)
and $T_{ab}^{(\phi)}$ then assumes the well-known dissipative fluid form
$T_{ab}=\rho u_{a}u_{b}+Ph_{ab}+\pi_{ab}+q_{a}u_{b}+q_{b}u_{a}\,,$ (5)
where $\rho$, $P$, $\pi_{ab}$ and $q_{a}$ are, respectively, an effective energy density, effective isotropic pressure, effective anisotropic stress tensor, effective heat flux density, while $h_{ab}\equiv g_{ab}+u_{a}u_{b}$ is the 3-metric on the 3-space seen by timelike observers with 4-velocity $u^{a}$. The effective fluid quantities associated with $T^{(\phi)}_{ab}$, i.e., $\rho^{(\phi)}$, $P^{(\phi)}$, $\pi_{ab}^{(\phi)}$ and $q_{a}^{(\phi)}$, as well as its kinematic quantities [25], are reported in [26] for first-generation ST gravity and in [27, 21] for its Horndeski generalization. Eckart [28] developed a theory of relativistic dissipative fluids which, although non-causal (the heat flux density $q^{a}$ is spacelike) and subject to instabilities, is the first step in the description of dissipative fluids and is much simpler than its relativistic generalizations [29, 30]. It is based on three assumptions in the form of constitutive relations for the fluid material, the most important being the generalization of Fourier’s law for heat conduction [28]
$q_{a}=-{\cal K}h_{ab}\left(\nabla^{b}{\cal T}+{\cal T}\dot{u}^{b}\right)\,,$ (6)
where ${\cal K}$ is the fluid’s thermal conductivity, ${\cal T}$ is its temperature, and $\dot{u}^{a}\equiv u^{b}\nabla_{b}u^{a}$ is its 4-acceleration. A priori there is no reason for the effective $\phi$-fluid of ST gravity to obey the Eckart constitutive relation (6) but, miraculously, it does: a direct computation of the effective fluid quantities and the comparison of $q_{a}^{(\phi)}$ and $\dot{u}^{a}$ yield [26, 19, 20, 21] $q_{a}^{(\phi)}=-{\cal KT}\dot{u}_{a}$, from which one identifies the product of effective thermal conductivity and effective temperature
${\cal KT}=\frac{\sqrt{-\nabla^{c}\phi\nabla_{c}\phi}}{8\pi\phi}\,.$ (7)
One cannot separate ${\cal K}$ and ${\cal T}$ and this is the closest that one comes to defining a “temperature of gravity”, which is relative to GR. GR, corresponding to ϕ=italic-$\phi=$ const. and ${\cal KT}=0$, is the “zero temperature” state, while the excitation of a non-trivial (i.e., propagating) scalar degree of freedom $\phi$ corresponds to a thermally excited state at ${\cal KT}>0$. The approach of ST gravity to GR (i.e., ${\cal KT}\to 0$), or its departure from it (i.e., increasing ${\cal KT}$), are described by the equation [26, 19, 20, 21, 24]
$\frac{d\left({\cal KT}\right)}{d\tau}=8\pi\left({\cal KT}\right)^{2}-\Theta{\cal KT}+\frac{\Box\phi}{8\pi\phi}\,,$ (8)
where $\tau$ is the proper time along the lines of the effective $\phi$-fluid and $\Theta\equiv\nabla_{c}u^{c}$ is its expansion scalar. Substituting the evolution equation of $\phi$ (3) gives
$\frac{d\left({\cal KT}\right)}{d\tau}=8\pi\left({\cal KT}\right)^{2}-\Theta{\cal KT}+\frac{T^{\mathrm{(m)}}}{\left(2\omega+3\right)\phi}+\frac{1}{8\pi\left(2\omega+3\right)}\left(V^{\prime}-\frac{2V}{\phi}-\frac{\omega^{\prime}}{\phi}\,\nabla^{c}\phi\nabla_{c}\phi\right)\,.$ (9)
Gravity is “heated” by positive terms in the right-hand side, which contribute to $d\left({\cal KT}\right)/d\tau>0$ causing departure from GR, and is “cooled” by negative ones contributing to $d\left({\cal KT}\right)/d\tau<0$ and to gravity approaching GR.
Since $\phi>0$ to keep the effective gravitational coupling $G_{\mathrm{eff}}$ positive, the matter contribution has the sign of $T^{\mathrm{(m)}}$.
Here we consider Brans-Dicke gravity [12] with $\omega=$ const., $V(\phi)\equiv 0$, for which Eq. (9) reduces to
$\frac{d\left({\cal KT}\right)}{d\tau}={\cal KT}\left(8\pi{\cal KT}-\Theta\right)+\frac{T^{\mathrm{(m)}}}{\left(2\omega+3\right)\phi}\,.$ (10)
The thermal view of ST gravity is only defined when $\nabla^{a}\phi$ is timelike and future-oriented and is not applicable when $\nabla^{a}\phi$ and $u^{a}$ are spacelike or null. In first-generation ST gravity, black hole no-hair theorems guarantee that, under suitable conditions, the spacetime geometry outside the horizon of a stationary, asymptotically flat black hole is the same as in GR. The no-hair theorems started with Hawking’s celebrated theorem establishing that stationary black holes that are asymptotically flat in vacuum Brans-Dicke gravity with a massless gravitational scalar field $\phi$ are the same as in GR [31]. Furthermore, the end state of gravitational collapse will produce stationary black holes, which are forced to be axisymmetric [32]. Over the years, the theorem was generalized to include spherical black holes in more general theories [33, 34, 35, 36] and realistic axisymmetric black holes in ST gravities in which the scalar field potential $V(\phi)$ has a minimum [37]. Likewise, the violation of no-hair and the search for black hole hair in alternative theories of gravity is the subject of a large literature (see the recent review [38]).
In theories descrived by the action (1), provided that $V(\phi)$ has a minimum at $\phi_{0}$ and that $\phi$ and $\omega(\phi)$ are not singular outside, or on, the horizon a vacuum static black hole in ST gravity is the same as in GR, i.e., a Kerr black hole because the scalar field is forced to have the constant value $\phi_{0}$ outside the horizon and then gravity reduces to GR. Static, asymptotically flat, spherical black holes in vacuum Brans-Dicke gravity are forced to be Schwarzschild, with line element
$ds^{2}=-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{1-2m/r}+r^{2}d\Omega_{(2)}^{2}\,,$ (11)
where $d\Omega_{(2)}^{2}\equiv d\vartheta^{2}+\sin^{2}\vartheta\,d\varphi^{2}$ is the line element on the unit 2-sphere and $m$ is the black hole mass.
What about the interior of black holes? Here we focus on spherical, asymptotically flat black holes. Even though the exterior geometry is static Schwarzschild, the interior is dynamic. The coordinates $t$ and $r$ switch roles at the horizon $r=2m$ and the interior region is dynamic, with time-dependent scalar field $\phi$ and there is a chance to apply the thermal view of ST gravity. We focus on the region near the black hole singularity, which is spacelike in Schwarzschild [18, 39]. We do not consider rotating or electrically charged black holes, for which the singularity can be timelike or null (possibly including Cauchy horizons) [18, 39]. It is remarkable that spacelike singularities in GR exhibit a high degree of universality. In seminal work on the approach to spacelike singularities, Belinski, Khalatnikov and Lifshitz [40, 41] have shown that spacelike singularities in GR are ultralocal and oscillatory. In the ultra-local limit, the vacuum geometry near a spacelike singularity is characterized by an infinite sequence of Kasner epochs, in each one of which the geometry is described by a Kasner solution of vacuum (i.e., $R_{ab}=0$) GR, with line element
$ds^{2}=-dt^{2}+a_{1}^{2}(t)dx^{2}+a_{2}^{2}(t)dy^{2}+a_{3}^{2}(t)dz^{2}$ (12)
in Cartesian coordinates, where
$a_{i}(t)=\left(\frac{t}{t_{0}}\right)^{p_{i}}\quad\quad(i=1,2,3)$ (13)
with constant $t_{0}$. The constant exponents $p_{i}$ satisfy
$p_{1}+p_{2}+p_{3}=1\,,$ (14)
$p_{1}^{2}+p_{2}^{2}+p_{3}^{2}=1\,.$ (15)
A unit volume of 3-space in comoving coordinates scales as $a^{3}(t)\equiv a_{1}(t)a_{2}(t)a_{3}(t)=t/t_{0}$. The Kasner solution is a special case of the spatially homogeneous and anisotropic Bianchi I geometry which generalizes the spatially flat Friedmann-Lemaître-Robertson-Walker universe. In the approach to a generic spacelike singularity, the exponents $p_{i}$ depend on the spatial coordinates, but this dependence disappears in the ultralocal limit. This phenomenology occurs also with certain forms of matter in GR, e.g., for perfect fluids with equation of state $P=w\rho$ and $w\leq 1/3$ [42], for which the singularity is dominated by gravity and “matter doesn’t matter”, but not for other forms of mass-energy (e.g., for perfect fluids with $P>\rho/3$ [42]).
Let us come to alternative theories of gravity. There are strong indications that the Kasner phenomenon persists in modified gravity. The Kasner exponents are, in general, functions of the spatial position and the geometry reduces to Kasner in the ultralocal limit. In higher-order gravity, where multiple energy scales are present due to multiple coupling constants in the action, sequences of “Kasner eons” have been discovered, each one of them consisting of a Kasner epoch with different Kasner exponents, with very brief transition periods between one epoch and the next. In particular, [43, 44, 45] investigate in this regard fourth order gravity, which contains $f(R)$ gravity as a special case. This class of theories is equivalent to an $\omega=0$ Brans-Dicke theory with $\phi=f^{\prime}(R)$ and a complicated potential [9, 10, 11] and it is not far-fetched to expect the Kasner phenomenon discovered there [43, 44, 46] to extend to at least first-generation ST gravity with second order equations of motion, and possibly to Horndeski [17] and DHOST gravity. Interesting behaviour consisting of sequences of Kasner eons has been found recently also in Lovelock gravity black holes [47, 46]. In first-generation ST gravity, where the exterior (static, asymptotically flat) black hole geometry is Schwarzschild, as guaranteed by the no-hair theorems, it is reasonable to expect the region near the spacelike black hole singularity to have Kasner geometry. Let us apply the thermal view to this region of ST black hole interiors.
Black hole interiors via the ST Kasner solution—The Kasner behaviour near the spacelike singularities of ST black holes is described by the Kasner solution of ST gravity [48, 42], which differs from the GR Kasner solution as follows [48, 42]. For definiteness, in this section we restrict to vacuum Brans-Dicke gravity described by the action (1) with constant $\omega$ and $V(\phi)=0$, for which the Kasner solution is well known, contrary to more general theories. The line element is given by Eqs. (12), (13), while $\phi=\phi(t)$ because of spatial homogeneity. In terms of the average scale factor $a(t)\equiv\Big{[}a_{1}(t)a_{2}(t)a_{3}(t)\Big{]}^{1/3}$ and of the individual directional Hubble functions $H_{i}(t)\equiv\dot{a}_{i}/a_{i}$, the field equations for the Kasner geometry (12) reduce to [48, 42]
$\frac{1}{a^{3}}\,\frac{d}{dt}\left(a^{3}H_{i}\right)+H_{i}\,\frac{\dot{\phi}}{\phi}=0\,,$ (16)
$\frac{d}{dt}\left(a^{3}\dot{\phi}\right)=0\,,$ (17)
$H_{1}H_{2}+H_{2}H_{3}+H_{1}H_{3}=\frac{\omega}{2}\left(\frac{\dot{\phi}}{\phi}\right)^{2}+\frac{\ddot{\phi}}{\phi}\,,$ (18)
where an overdot denotes differentiation with respect to the comoving time $t$. They admit the first integrals
$\dot{\phi}=\frac{C_{1}}{a^{3}}\,,$ (19)
$\frac{d(a^{3})}{dt}=\frac{C_{2}}{\phi}\,,$ (20)
with $C_{1,2}$ integration constants. The solution is [48, 42]
$a_{i}(t)$ $=\left(\frac{t}{t_{0}}\right)^{\frac{p_{i}}{1+C}}\,,$ (21)
$\phi(t)$ $=$ $\phi_{0}\left(\frac{t}{t_{0}}\right)^{\frac{C}{1+C}}\,,\quad\quad C=\frac{C_{1}}{C_{2}}\,.$ (22)
Again, $p_{1}+p_{2}+p_{3}=1$ but now $a^{3}=\left(t/t_{0}\right)^{\frac{1}{C+1}}$ and [48, 42]
$p_{1}^{2}+p_{2}^{2}+p_{3}^{2}=1-C\left(\omega C-2\right)\,,$ (23)
which marks the difference with GR. The ST Kasner solution reduces to the GR one for $C=0$. It was already noted in [42] that the scalar degree of freedom alters substantially the GR behavior at the initial singularity. The thermal view interprets this situation as the divergence of gravity from GR, with ${\cal KT}$ going infinitely far away from zero.
In order for the timelike vector field $\nabla^{a}\phi$ to be future-oriented it must be
$\dot{\phi}=\frac{C}{C+1}\,\frac{\phi_{0}}{t_{0}}\left(\frac{t_{0}}{t}\right)^{\frac{1}{C+1}}<0\,;$ (24)
since $\phi_{0}>0$, it must be $\frac{C}{C+1}<0$, which is satisfied in the range
$-1<C<0\,.$ (25)
Outside of this range of values of $C$, the thermal view of ST gravity is not applicable. For these values of $C$, the 3-volume $a^{3}\to 0$, while $\phi\to+\infty$ and $G_{\mathrm{eff}}\to 0$ at the singularity $t\to 0^{+}$.
To quantify the departure from GR near the spacelike singularity, note that
${\cal KT}=\frac{|\dot{\phi}|}{8\pi\phi}=\frac{|C|}{8\pi\left(C+1\right)t}\,,$ (26)
diverges as $1/t$ at the singularity $t\to 0^{+}$. By contrast, the GR Kasner solution corresponding to $C=0$ has constant $\phi$ and ${\cal KT}=0$ and the time-dependence of $a_{i}(t)$ and $\phi(t)$, although still power-law, is quite different from the GR case, as noted in [48, 42]. Hence, in the absence of matter and scalar field potential, Brans-Dicke gravity diverges away from GR near spacelike singularities, confirming an insight proposed in [19, 20, 21]: gravity is “hot” near these singularities. Although the discussion is simple, the universality of the Kasner behaviour near spacelike singularities is quite powerful, allowing general conclusions about the departure of gravity from GR in these regions.
Ref. [42] also reports the special degenerate solution for $C=-1$ in the parameter range $-3/2<\omega<-4/3$:
$a_{i}(t)$ $=$ $a_{i}^{(0)}\,\mbox{e}^{p_{i}\,t/t_{0}}\,,$ (27)
$\phi(t)$ $=$ $\phi_{0}\,\mbox{e}^{-\,t/t_{0}}\,,$ (28)
for which $\dot{\phi}=-\phi/t_{0}<0$ and $\nabla^{a}\phi$ is future-oriented. This solution is reported also in [48] with the incorrect sign in the argument of the exponential in $\phi(t)$, but it is straightworward to check that (27) is correct using Eq. (19) with $C=-1$. For this solution,
${\cal KT}=\frac{|\dot{\phi}|}{8\pi\phi}=\frac{1}{8\pi t_{0}}$ (29)
and this solution always remains different from the ($C=0$) GR solution, but at a finite constant “distance” ${\cal KT}$ from it, which seems bizarre. To understand this situation, note that the expansion scalar
$\Theta=\frac{1}{a^{3}}\,\frac{d(a^{3})}{dt}=\frac{1}{t_{0}}=8\pi{\cal KT}$ (30)
and, since we are in vacuo,
$\frac{d\left({\cal KT}\right)}{dt}={\cal KT}\left(8\pi{\cal KT}-\Theta\right)=0\,,$ (31)
allowing ${\cal KT}$ to remain constant. Indeed, Ref. [49] identified the $\left(\Theta,{\cal KT}\right)$ plane as a convenient description of vacuum ST gravity and the line $8\pi{\cal KT}=\Theta$ as a line of fixed points that cannot be crossed dynamically. For expanding 3-spaces with $\Theta>0$, this line separates the region $8\pi{\cal KT}>\Theta$, in which $d\left({\cal KT}\right)/d\tau>0$ and gravity always departs from GR, from the other region $d\left({\cal KT}\right)/d\tau<0$ in which ST gravity always converges to GR [49]. The fine-tuned solution (27), (28) lies on this special line, along which $8\pi{\cal KT}=\Theta=$ const.
Including matter, $V(\phi)$ and $\omega(\phi)$—Let us focus now on Eq. (10) including matter, where the proper time $\tau$ along the effective fluid lines coincides with the Kasner comoving time $t$ because
$u^{0}\equiv\frac{g^{0b}\partial_{b}\phi}{\sqrt{-\nabla^{c}\phi\nabla_{c}\phi}}=\frac{-\dot{\phi}}{|\dot{\phi}|}=1=\frac{dt}{d\tau}\,.$ (32)
Furthermore, assume that matter is described by a perfect or imperfect fluid with linear barotropic equation of state $P=w\rho$ with $w=$ const., then $T^{\mathrm{(m)}}=-\rho+3P=\left(3w-1\right)\rho$. A radiation fluid with $P=\rho/3$ (or any form of conformal matter, which has $T^{\mathrm{(m)}}=0$) does not affect ${\cal KT}$. An imperfect (real) fluid has the form (5) and, as for perfect fluids, its trace is $T^{\mathrm{(m)}}=-\rho+3P$.
Let us estimate the powers with which the three terms in the right-hand side of Eq. (10) scale as $t\to 0^{+}$. Following the authors of [42], who discuss the field equations instead of Eq. (10), we use the vacuum solution to estimate the powers in the gravitational degrees of freedom. This procedure is justified by the fact that the fluid density and pressure scale the same way with $a$ for a test fluid or for a gravitating one and is corroborated by the exact Bianchi I solutions for gravitating matter [42].
In the ST vacuum Kasner solution we have ${\cal KT}\sim 1/t$, $\Theta\sim 1/t$, while $\rho\sim P=w\rho\sim a^{-3(w+1)}\sim t^{-\,\frac{w+1}{C+1}}$, therefore $T^{\mathrm{(m)}}/\phi\sim 1/t^{\frac{w+C+1}{C+1}}$. The gravity terms $\left({\cal KT}\right)^{2}\sim t^{-2}$ and  $-\Theta{\cal KT}\sim t^{-2}$ dominate over the matter term $\frac{T^{\mathrm{(m)}}}{\left(2\omega+3\right)\phi}$ as $t\to 0^{+}$ if
$\frac{w+C+1}{C+1}<2$ (33)
which (since $C+1>0$) is equivalent to $w<C+1$. On the contrary, the matter term
$\frac{T^{\mathrm{(m)}}}{\left(2\omega+3\right)\phi}=\frac{(3w-1)\rho_{0}\,t_{0}^{\frac{w+C+1}{C+1}}}{\left(2\omega+3\right)\phi_{0}\,t^{\frac{w+C+1}{C+1}}}$
dominates if $w>C+1$. The same conclusion is reached in [42] by comparing terms in the field equations and using the vacuum Kasner solution to estimate the gravitational and matter terms (cf. Eq. (15) of [42]).
Gravity and the matter term scale with the same power of $1/t$ as $t\to 0^{+}$ if $w=C+1>0$, in which case (not discussed in [42] or subsequent literature) we have
$\frac{d\left({\cal KT}\right)}{dt}=\left[\frac{C}{8\pi(C+1)}+\frac{\left(3C+2\right)\rho_{0}t_{0}^{2}}{\phi_{0}}\right]\frac{1}{t^{2}}\,;$ (34)
since $-1<3C+2<2$, the sign of the right-hand side depends on the initial conditions, which determine $\rho_{0},\phi_{0},t_{0}$, and on $w=C+1$. If $-1<C<-2/3$, one can in principle tune the initial condition so that ${\cal KT}=$ const. (similarly to the line $8\pi{\cal KT}=\Theta$ for vacuum), although such fine-tuning would be physically uninteresting.
In addition to matter, one can think of including in the picture a scalar field potential $V(\phi)$, which complicates the analysis significantly. If $V$ is quadratic, it disappears from Eq. (3) for $\Box\phi$ and from Eq. (10) for ${\cal KT}$ because it only enters these equations through the combination $\phi V^{\prime}-2V$. Otherwise, $V(\phi)$ makes Eq. (3) nonlinear and difficult to solve. As noted in [49] using Eq. (9), if $V(\phi)$ grows with $\phi$ faster than $\phi^{2}$, it “heats up” gravity, while it “cools” it if it grows slower than $\phi^{2}$. Likewise, if present, the last term in Eq. (9), $\frac{\omega^{\prime}\dot{\phi}^{2}}{8\pi\left(2\omega+3\right)\phi}$, contributes to “heating gravity” if $\omega(\phi)$ is monotonically increasing. However, without an explicit solution one cannot estimate the size of the terms in the right-hand side of Eq. (10) and decide which ones are dominant.
Conclusions—We have extended the thermal view of ST gravity to cover the interior region of ST black holes near their spacelike singularities, thus filling an obvious gap in the applications of this new formalism. The discussion of the previous section appears simple, perhaps too simple, but it is a virtue of the new thermal view of ST gravity that it simplifies the description of ST physics and offers a unified frame to interpret technical results in a coherent way. The universality of the Kasner behaviour near spacelike singularities allows one to reach quite general conclusions about the departure of gravity from GR in these regions, where the spacetime structure fails. A first implication is that, if quantization adds a scalar degree of freedom to gravity, as in $f(R)$ gravity [9, 10, 11] or in Starobinski inflation [5], then one should not attempt to quantize the Einstein when looking for Kasner transitions in singularity-free bouncing universes (which are of great current interest), but one should address the field equations of ST gravity instead [50], or adopt an approach not tied to the classical field equations. The conclusion may even be more radical: if the thermal view is taken seriously, gravity and its field equations could be emergent and not suitable for quantization, as proposed in Jacobson’s thermodynamics of spacetime [51, 52]. The latter is quite similar in spirit, although completely different in substance, to the thermal view of ST gravity.
We have not said anything about timelike singularities such as those encountered in electrically charged black holes [18, 39] or about null singularities (such as thunderbolt singularities [53]) or Cauchy horizons. Although the general idea that gravity is “hot” (i.e., deviates from GR) near all spacetime singularities presumably remains valid [19, 20, 21], no generic behaviour like the Kasner phenomenology is known near these other types of singularities. Explicit examples aiming to falsify or support the general idea that such spacetime singularities also imply a catastrophic divergence from GR will be sought for elsewhere.
This work is supported, in part, by the Natural Sciences & Engineering Research Council of Canada (grant 2023-03234).

### v2

@@
 Over the years, the theorem was generalized to include spherical black holes in more general theories [33, 34, 35, 36] and realistic axisymmetric black holes in ST gravities in which the scalar field potential $V(\phi)$ has a minimum [37].
-Likewise, the violation of no-hair and the search for black hole hair in alternative theories of gravity is the subject of a large literature (see the recent review [38]).
-In theories descrived by the action (1), provided that $V(\phi)$ has a minimum at $\phi_{0}$ and that $\phi$ and $\omega(\phi)$ are not singular outside, or on, the horizon a vacuum static black hole in ST gravity is the same as in GR, i.e., a Kerr black hole because the scalar field is forced to have the constant value $\phi_{0}$ outside the horizon and then gravity reduces to GR.
+Likewise, the violation of no-hair theorems and the search for black hole hair in alternative theories of gravity is the subject of a large literature (see the recent review [38]).
+In theories descrived by the action (1), provided that $V(\phi)$ has a minimum at $\phi_{0}$ and that $\phi$ and $\omega(\phi)$ are not singular outside, or on, the horizon a vacuum stationary black hole in ST gravity is the same as in GR, i.e., a Kerr black hole because the scalar field is forced to have the constant value $\phi_{0}$ outside the horizon and then gravity reduces to GR.
 Static, asymptotically flat, spherical black holes in vacuum Brans-Dicke gravity are forced to be Schwarzschild, with line element
@@
 The universality of the Kasner behaviour near spacelike singularities allows one to reach quite general conclusions about the departure of gravity from GR in these regions, where the spacetime structure fails.
-A first implication is that, if quantization adds a scalar degree of freedom to gravity, as in $f(R)$ gravity [9, 10, 11] or in Starobinski inflation [5], then one should not attempt to quantize the Einstein when looking for Kasner transitions in singularity-free bouncing universes (which are of great current interest), but one should address the field equations of ST gravity instead [50], or adopt an approach not tied to the classical field equations.
-The conclusion may even be more radical: if the thermal view is taken seriously, gravity and its field equations could be emergent and not suitable for quantization, as proposed in Jacobson’s thermodynamics of spacetime [51, 52].
+A first implication is that, if quantization adds a scalar degree of freedom to gravity, as in $f(R)$ gravity [9, 10, 11] or in Starobinski inflation [5], then one should not attempt to quantize the Einstein equations when looking for Kasner transitions in singularity-free bouncing universes (which are of great current interest), but one should address the field equations of ST gravity instead [50], or adopt an approach not tied to the classical field equations.
+The conclusion may even be more radical: if the thermal view is taken seriously, gravity and its field equations could be emergent and not suitable for quantization as proposed, e.g., in Jacobson’s thermodynamics of spacetime [51, 52].
 The latter is quite similar in spirit, although completely different in substance, to the thermal view of ST gravity.
@@
 Explicit examples aiming to falsify or support the general idea that such spacetime singularities also imply a catastrophic divergence from GR will be sought for elsewhere.
+Singularities are still present in scalar-tensor gravity and the thermal formalism, being purely classical, does nothing to cure them.
+All that it offers is the realization of how these spacetime singularities are different from those in GR and why—a cure for these singularities will have to come from somewhere else.
+The understanding of “different” singularities proposed here may have some impact on certain approaches to quantization, but there is no reason to believe this impact to be crucial.
+The main point of the new thermal view of scalar-tensor gravity is that it offers a new framework to understand apparently unrelated phenomena that have appeared in the literature over three decades.
+They are indeed related and the new thermal formalism, which is rather concise, unifies apparently disconnected phenomena in a single view, simplifying their description.
+However, all these aspects of scalar-tensor gravity need to be revisited one by one in the new picture and then the unifying character of the latter is not manifest.
+We refer the reader to [19, 20, 21, 22, 23, 24] for details and we will extend the study to new phenomena in future work.
 This work is supported, in part, by the Natural Sciences & Engineering Research Council of Canada (grant 2023-03234).